OpenStudy (anonymous):

I know this should be simple, but I still got it wrong... Another related rates, implicit differentiation problem: The base x of the right triangle below increases at a rate of 5 cm/s, while the height remains constant at h = 20. How fast is the angle θ changing when x = 20? I'll post diagrams and formulas for what I attempted to setup.

OpenStudy (anonymous):

\[\tan \theta=20/x\] then implicitly differentiated.

OpenStudy (anonymous):

I get \[\sec^2(\theta)*d \theta/dt =-20(dx/dt)/x^2\] solve for theta rate of change

OpenStudy (anonymous):

I got \[-20(dx/dt)\cos^2\theta/x^2\] because 1/sec^2 is cos^2? I hope.... am I starting completely wrong? have I missed a sign somewhere?

OpenStudy (anonymous):

 Am I assuming the right angle \(\theta\)?

OpenStudy (anonymous):

yes, sorry.... forgot to post that.

OpenStudy (anonymous):

I think you're doing great so far. All you need now is to substitute in the equation you've got for the rate of change.

OpenStudy (anonymous):

You're right I found it.... I just multiplied wrong... I couldn't figure out why it wasn't coming out right and I doubted I had started right because of the sec^2 to cos^2.... yes, I have it now.

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

thank you... now to the last two problems... see if I can solve them myself first...